### v1

# from Monocular RGB-D Video

Abstract
[…]
1 Introduction
The challenge of reconstructing the dense 3D geometry, appearance, and motion of dynamic scenes from videos has persisted for decades, offering a wide range of applications in virtual reality and augmented reality. Due to the inherent ambiguities of the problem, traditional dense non-rigid reconstruction approaches often assume strong priors on the object’s shape and motion. Model-based approaches [75, 1, 26, 11, 65, 25] require 3D parametric models [2, 36, 42, 6, 28, 84] of specific object categories such as humans [2, 36], faces [6, 28], or animals [84]. While these methods achieve promising reconstruction results, they fall short of generalizing to arbitrary object categories. In this paper, we focus on model-agnostic approaches that could generalize to generic non-rigid objects. This task is often approached by jointly fusing a canonical shape and a per-frame deformation field from an RGB-D video sequence [41, 24, 59, 60, 8, 7]. Due to the inherent constraints of physical sensors and the use of traditional voxel-grid representations, they often struggle to produce high-fidelity surface details and can only reconstruct observed regions.
Recent attention has shifted towards leveraging neural representations for scene reconstruction. Neural Radiance Fields (NeRF) [39] represent scene density and color in the weights of a neural network. In combination with volume rendering, NeRF achieves unprecedented performance on novel view synthesis. Many follow-up works adapt this idea into surface reconstruction [79, 72, 4, 71], SLAM [61, 82, 27], dynamic scene reconstruction [48, 44, 45, 9] and more. Notably, NDR [9] proposes the first RGB-D neural implicit dynamic surface reconstruction method. Thanks to the expressiveness of neural representations, they achieve accurate and smooth surface reconstruction. However, real-world captures often contain large unobserved regions. The inherent smoothness of neural representations, and the difficulty to incorporate domain-specific priors, causes current methods to fall short of achieving realistic completion.
[…]
Our key contribution is to demonstrate the capability to learn metrically accurate geometry and deformations of dynamic objects from casually captured RGB-D videos while achieving realistic completion in unobserved regions with diffusion priors. We propose a temporal view-dependent SDS strategy to achieve photo-realistic completion while learning an accurate deformation field. A canonical space regularization strategy is used to avoid the trivial solution of surface completion being explained via deformations. Experimental results across various datasets show that MorpheuS achieves competitive reconstruction accuracy and realistic completion.
2 Related Work
Dense Non-rigid Reconstruction. Traditional dense dynamic scene reconstruction methods often require strong prior knowledge about the object shape and motion, fitting 3D prior parametric models [36, 51, 84, 28, 75, 1, 26, 11, 65, 25] to video sequences of a specific object category, and thus cannot work with arbitrary unseen objects. Early model-agnostic methods either require accurate dense correspondences [19] or a template shape of the object to be reconstructed first before tracking the dynamic motion [62, 83, 80, 5]. The seminal work DynamicFusion [41] was the first direct and template-free method to achieve real-time simultaneous tracking and reconstruction of generic non-rigid objects via joint optimization of a canonical shape and a per-frame deformation field. Follow-up methods improve the robustness of motion tracking and reconstruction quality by incorporating photo-metric constraints [20], classical  [24] or learned correspondences [8, 7], reformulating the optimization as direct dense SDF alignment [59, 60]. Although these methods have shown promising reconstruction results, they fall short of completing the unobserved regions due to their inherent traditional scene representation and lack of generative priors.
Neural Rendering for Dynamic Scenes. The success of NeRF [39] in novel view synthesis has inspired many follow-up works [48, 44, 29, 45, 17, 66, 68, 15, 9, 18, 31] in dynamic scene reconstruction. These approaches can be broadly classified into two paradigms: a) Directly modeling the scene geometry and motion as a unified 4D space-time field in the world (observation) space [29, 17, 10, 16, 31, 54, 23], where motion priors, such as scene flow or trajectory can be involved as an additional regularization [29, 68, 31]. b) Learning a canonical shape by warping the shape from observation space to canonical space via a separate deformation field [48, 44, 45, 66, 15, 9, 18, 76, 14]. Among these works, NDR [9] is particularly relevant to our work. They achieve high-fidelity dynamic surface reconstruction from monocular RGB-D inputs. However, NDR falls short of realistic reconstruction in unobserved regions.
Neural Rendering with Diffusion Priors. Recent years have witnessed remarkable progress in diffusion-based foundation models [50, 52], unlocking unprecedented capabilities for diverse applications such as image synthesis, editing, and shape generation [46]. DreamFusion [47] is a notable example that leverages pre-trained image diffusion priors to optimize a NeRF model through SDS, achieving high-quality text-to-3D generation. Building upon DreamFusion’s success, numerous follow-up works have emerged, particularly in text-to-3D [32, 12, 38, 69, 74, 81], image-to-3D [37, 77, 49, 64], scene editing [22, 53, 43], etc. Some other works [34, 56, 35] try to enhance the generation quality via fine-tuning diffusion prior. A noteworthy example is Zero-1-to-3 [34], which fine-tunes the stable-diffusion [50] model with relative camera pose conditioning. The resulting model takes an input image and the camera parameters of the viewpoint, and outputs an image rendered at the novel view. Recently, several works have been proposed for dynamic scene generation [58] and editing [53, 43]. MAV3D [58] adopts a text-to-video diffusion model [57] for text-to-4d generation. Control4D [53] utilizes a 2D diffusion prior to train a 4D GAN, enabling 4D editing of human portraits. AvatarStudio [43] fine-tunes the diffusion prior with time stamps and viewpoints, enabling editing of human head avatars with a view-and-time aware SDS loss. Different from these works, our focus is learning metrically accurate motion and geometry from the input monocular RGB-D video observations while achieving realistic completion in the unobserved region with a diffusion prior.
Figure 2: Overview of MorpheuS. 1) Dynamic surface rendering: we model the target dynamic scene via a deformation field that maps a point from observation space to a hyper-dimensional canonical space and a canonical field that decodes the point into SDF and color. 2) Diffusion prior: we leverage a diffusion prior and perform SDS to complete the unobserved region. Note here the de-noising process is in latent space. All visualization is generated via decoding the latent vector for illustration purposes. 3) Optimization: We optimize the scene representation using real view supervision ${\mathcal{L}}_{\mathrm{real}}$, SDS loss ${\mathcal{L}}_{\mathrm{S}}$, canonical regularization ${\mathcal{L}}_{\mathrm{reg}}^{\mathrm{cano}}$, and parameter regularization ${\mathcal{L}}_{\mathrm{reg}}^{\mathrm{param}}$.
3 Method
[…]
3.1 Dynamic Neural Surface Rendering
Deformation Field. Inspired by the recent success of dynamic scene reconstruction [9], we model the deformation of the scene via a deformation network $D(\cdot)$ and a topology network $T(\cdot)$. Given a point ${\mathbf{x}}_{t}\in\mathbb{R}^{3}$ in observation space at frame $t$, $D(\cdot)$ and $T(\cdot)$ map the point from the observation space to a point in the hyper-dimensional canonical space ${\mathbf{x}}^{\prime}\in\mathbb{R}^{3+m}$:
${\mathbf{x}}^{\prime}=\{{\mathbf{x}}^{\prime}_{m},{\mathbf{x}}^{\prime}_{a}\}=\{{\mathbf{x}}_{t}+D(\phi({\mathbf{x}}_{t}),{\mathcal{V}}_{t}(t)),T(\phi({\mathbf{x}}_{t}),{\mathcal{V}}_{t}(t))\},$ (1)
where ${\mathbf{x}}^{\prime}_{m}\in\mathbb{R}^{3}$ is the 3D coordinate of the deformed point and ${\mathbf{x}}^{\prime}_{a}\in\mathbb{R}^{m}$ is the ambient coordinate predicted by the topology network $T(\cdot)$. $\phi(\cdot)$ and ${\mathcal{V}}_{t}(\cdot)$ are encoding functions of the location ${\mathbf{x}}_{t}$ and time $t$. We use a frequency encoding $\phi(\cdot)$ to model the position ${\mathbf{x}}_{t}$ for simplicity. For ${\mathcal{V}}_{t}(\cdot)$, we use a multi-resolution 1-D feature grid instead of a per-frame deformation code [45, 9]. The feature vector for each frame $t$ is obtained via linear interpolation.
Hyper-dimensional Canonical Field. The 3D geometry and appearance of the target object are represented as the zero-level set of the signed distance function (SDF) field $s$ and color ${\mathbf{c}}$:
$s,{\mathbf{h}}=f_{\gamma}({\mathcal{V}}_{s}({\mathbf{x}}^{\prime}_{m}),{\mathbf{x}}^{\prime}_{a}),$ (2)
${\mathbf{c}}=f_{\alpha}({\mathcal{V}}_{c}({\mathbf{x}}^{\prime}_{m}),{\mathbf{h}}),$ (3)
where ${\mathcal{V}}_{s}$ and ${\mathcal{V}}_{c}$ represent the encoding of the input position ${\mathbf{x}}^{\prime}_{m}$. Here we choose the Hash encoding [40] for the fast optimization speed. To incorporate the diffusion prior, a good initialization is important for the canonical field. Existing works in text-to-3D tasks often choose to initialize the radiance field with a fixed blob function [47, 32]. However, we find that the deformation field fails to converge with the fixed blob function as an initialization. Thus, we adopt the geometric initialization following SAL [3]. Since it is not straightforward to map the Hash features with ambient coordinates into a sphere, we append the 3D coordinate ${\mathbf{x}}^{\prime}_{m}$ and use zero initialization to effectively mask out the remaining coordinates in the input layer of the decoder $f_{\gamma}(\cdot)$.
Volume Rendering. We transform the SDF into volume density $\sigma_{\beta}(s)$ following VolSDF [79], where $\beta$ is a learn-able parameter which controls the smoothness of the surface. The volume density is then used for color $C$ and depth $D$ rendering:
$\begin{gathered}C=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},\quad D=\sum_{i=1}^{N}T_{i}\alpha_{i}d_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),\end{gathered}$ (4)
where $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$, $\delta_{i}$ is the distance between neighboring sample points.
[…]
We incorporate Zero-1-to-3 [34], a view-conditioned latent diffusion model, and distill knowledge from it to complete the unobserved region using SDS [47]:
$\nabla_{\theta}{\mathcal{L}}_{\mathrm{S}}={\mathbb{E}}_{i,\epsilon}\left[w(i)\left(\epsilon_{\psi}\left(\tilde{{\mathbf{z}}}_{i};{\mathbf{I}}_{r},\Delta{\mathbf{P}},i\right)-\epsilon\right)\frac{\partial{\mathbf{I}}_{v}}{\partial\theta}\right],$ (5)
where ${\mathbf{I}}_{v}$ is the rendered image from the sampled virtual view, which has a relative pose $\Delta{\mathbf{P}}$ with respect to the real image observation ${\mathbf{I}}_{r}$ at the reference view. Here the relative pose $\Delta{\mathbf{P}}=[\Delta r,\Delta\omega,\Delta\phi]=[r_{v}-r_{r},\omega_{v}-\omega_{r},\varphi_{v}-\varphi_{r}]$ is parameterised in polar-coordinate. Conditioned on the reference view ${\mathbf{I}}_{r}$, the relative pose $\Delta{\mathbf{P}}$, and the sampled time-step $i\sim\mathcal{U}(0,1)$, the de-noising U-Net $\epsilon_{\psi}(\cdot)$ applies a de-noising step on the noisy latent $\tilde{{\mathbf{z}}}_{i}$ of the rendered view ${\mathbf{I}}_{v}$. $\theta$ is the learn-able parameters of our dynamic scene model. $w(i)$ is a time-dependent weighting term.
Temporal View-dependent SDS. As our input is a monocular video of a dynamic object, we condition the diffusion model on different frames so as to capture the motion of the target object. We select a keyframe for every $k$ frame as the conditioning reference views of the diffusion prior. For any given rendered virtual view ${\mathbf{I}}_{v}^{t}$ at frame $t$, we choose its nearest keyframe $t_{k}$ as the reference view condition and modify the Eq. 5 as:
$\nabla_{\theta}{\mathcal{L}}_{\mathrm{S}}^{t}={\mathbb{E}}_{i,\epsilon}\left[\hat{w}(i)\left(\epsilon_{\psi}\left(\tilde{{\mathbf{z}}}_{i}^{t};{\mathbf{I}}_{r}^{t_{k}},\Delta{\mathbf{P}}^{t},i\right)-\epsilon\right)\frac{\partial{\mathbf{I}}_{v}^{t}}{\partial\theta}\right].$ (6)
The latent vectors of each keyframe for conditioning are pre-computed to save the computational cost. In practice, we find that using SDS does not have strict consistency across different frames, which affects the convergence of the deformation field. Thus, we incorporate a view-dependant modulation term to the weighting term $\hat{w}(i)=w(i)(\exp(|\Delta{\mathbf{P}}^{t}_{(1:2)}|)-1)$ to encourage the SDS loss to focus on novel views with larger view difference with respect to the reference view:
$|\Delta{\mathbf{P}}_{(1:2)}^{t}|=\arccos{\big{(}\frac{{\mathbf{o}}_{r}\cdot{\mathbf{o}}_{v}}{\left\lVert{\mathbf{o}}_{r}\right\rVert\left\lVert{\mathbf{o}}_{v}\right\rVert}\big{)}}/\pi,$ (7)
where ${\mathbf{o}}_{r}$ and ${\mathbf{o}}_{v}$ are the Cartesian coordinates of the camera frame origins of the reference and virtual views.
Implementation Details. Following Dreamfusion [47], we use the shading model for rendering virtual views and randomly replace albedo with white color to produce textureless rendering. A random background is used for the rendered image from a sampled virtual view. In order to ensure the convergence of the deformation field, we perform SDS every $j$ iteration. The SDS loss is generated via a reparameterization trick [63, 21]:
${\mathcal{L}}_{S}^{t}=\|\mathrm{stopgrad}({\mathbf{z}}_{i}^{t}-\mathrm{grad})-{\mathbf{z}}_{i}^{t}\|^{2},$ (8)
where $\mathrm{grad}=\hat{w}(i)\left(\epsilon_{\psi}\left(\tilde{{\mathbf{z}}}_{i}^{t};{\mathbf{I}}_{r}^{t_{k}},\Delta{\mathbf{P}}^{t},i\right)-\epsilon\right)$. ${\mathcal{L}}_{S}^{t}$ will be combined with a real view loss for optimization.
3.3 Optimization
The optimization of our scene representation can be formulated as:
${\mathcal{L}}={\mathcal{L}}_{\mathrm{real}}+{\mathcal{L}}_{\mathrm{S}}+{\mathcal{L}}_{\mathrm{reg}}^{\mathrm{cano}}+{\mathcal{L}}_{\mathrm{reg}}^{\mathrm{param}}.$ (9)
Real View Supervision. Given a sampled input frame $I_{t}$, we randomly sample a batch of pixels and perform volume rendering by casting rays through those pixels and transforming the ray points to the world coordinate with the corresponding camera pose. The real-view supervision is formulated as:
${\mathcal{L}}_{\mathrm{real}}=\underbrace{{\mathcal{L}}_{\mathrm{c}}+{\mathcal{L}}_{\mathrm{d}}+{\mathcal{L}}_{\mathrm{m}}}_{\text{rendering loss}}+\underbrace{{\mathcal{L}}_{\mathrm{sdf}}+{\mathcal{L}}_{\mathrm{surf}}+{\mathcal{L}}_{\mathrm{smooth}}}_{\text{per-point loss}}.$ (10)
We use $\ell_{2}$ loss for color ${\mathcal{L}}_{\mathrm{c}}$ and depth ${\mathcal{L}}_{\mathrm{d}}$, cross-entropy loss for mask ${\mathcal{L}}_{\mathrm{m}}$. In terms of the per-point loss, following [4, 71, 70] we set up a truncation distance $tr$ and apply pseudo-SDF loss ${\mathcal{L}}_{\mathrm{sdf}}$ to the ray-points within the truncation region. To speed up the convergence of our model, we employ ${\mathcal{L}}_{\mathrm{surf}}$ that directly supervises the SDF and color of the back-projected surface point using the input color and depth. Since Hash encoding does not have global continuity, resulting in noisy surface reconstruction, we further apply a normal smoothness term ${\mathcal{L}}_{\mathrm{smooth}}$ to near-surface points to encourage local smoothness of the SDF gradient in observation space:
${\mathcal{L}}_{\mathrm{smooth}}=\frac{1}{|S|}\sum_{{\mathbf{x}}\in S}\left\lVert\nabla s({\mathbf{x}})-\nabla s({\mathbf{x}}+\delta{{\mathbf{x}}})\right\rVert^{2},$ (11)
where $S$ is obtained by uniformly sampling $n$ points from $\hat{D}-tr/2$ to $\hat{D}+tr/2$ along sampled rays. The perturbed point ${\mathbf{x}}+\delta{\mathbf{x}}$ is sampled on a circle centered at ${\mathbf{x}}$ with radius $r$ orthogonal to the gradient vector $\nabla s({\mathbf{x}})$. Following [30] we compute the SDF gradient via finite difference.
Canonical Space Regularization. We find that performing regularization using all points on sampled rays in observation space leads to instability in learning the deformation of novel views. Instead, we propose to apply constraints directly to each individual slice of the hyper-dimensional canonical space by shortcutting the deformation network:
${\mathbf{x}}^{\prime}_{\mathrm{reg}}=\{{\mathbf{x}}_{t},T(\phi({\mathbf{x}}_{t}),{\mathcal{V}}_{t}(t))\}.$ (12)
[…]
${\mathcal{L}}^{\mathrm{cano}}_{\mathrm{reg}}={\mathcal{L}}_{\mathrm{ori}}+{\mathcal{L}}_{\mathrm{normal}}+{\mathcal{L}}_{\mathrm{eik}},$ (13)
where ${\mathcal{L}}_{\mathrm{ori}}$ is the orientation loss used in DreamFusion [47]. ${\mathcal{L}}_{\mathrm{eik}}$ is the eikonal loss that enforces the SDF gradient to have a unit norm. ${\mathcal{L}}_{\mathrm{normal}}$ is the normal smoothness regularization on each sampled point.
Parameter Space Regularization. We also perform regularization on the parameters of our model:
${\mathcal{L}}^{\mathrm{param}}_{\mathrm{reg}}={\mathcal{L}}_{\text{code}}+{\mathcal{L}}_{\beta}.$ (14)
${\mathcal{L}}_{\text{code}}$ is a regularization term on deformation code to encourage temporal smoothness of object motion:
${\mathcal{L}}_{\text{code}}=\frac{1}{|W_{t}|}\sum_{k\in W_{t}}\left\lVert 2{\mathcal{V}}_{t}(k)-{\mathcal{V}}_{t}(k-1)-{\mathcal{V}}_{t}(k+1)\right\rVert^{2},$ (15)
where $W_{t}=\{\dots,t-1,t,t+1,\dots\}$ is a local window around $t$. ${\mathcal{L}}_{\beta}$ is $\ell_{1}$ loss used for minimizing $\beta$ in $\sigma_{\beta}(s)$.
4 Experiments
4.1 Experimental Setup
Reference Frames NDR (Mesh) Ours (Mesh) NDR (Color) Ours (Color) Reference Frames NDR (Mesh) Ours (Mesh) NDR (Color) Ours (Color)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Real-world dataset reconstruction results (From left to right, top to bottom: Frog, Teddy, Human2, and Mochi). NDR [9] achieves high-quality surface reconstruction in the observed region but fails to produce photo-realistic completion, resulting in spurious surfaces in unobserved regions. In contrast, our method can produce high-quality 360° surface reconstruction.
Method Metric KillingFusion [59] DeepDeform [8] iPhone [18]
Avg.
Frog Duck Snoopy Human1 Human2 Human3 Teddy Mochi Haru
NDR Acc. [cm] $\downarrow$
1.25
1.34
1.32
0.74
0.83
0.83
6.12
2.19
2.73
1.92
Comp. [cm] $\downarrow$
0.99
1.15
1.05
0.63
0.68
0.65
1.91
1.06
1.73
1.09
Clip sim. $\uparrow$
82.86
82.52
80.66
82.22
83.15
79.24
75.83
74.12
77.19
79.75
Ours Acc. [cm] $\downarrow$ 0.77 0.88 1.05 0.57 0.61 0.78 0.99 0.77 1.51 0.88
Comp. [cm] $\downarrow$ 0.67 0.81 0.92 0.49 0.56 0.61 0.93 0.60 1.47 0.78
Clip sim. $\uparrow$ 90.24 93.15 90.82 86.41 88.75 82.67 80.50 81.57 86.87 86.77
Table 1: Per-scene quantitative results on real world datasets, including KillingFusion [59], DeepDeform [8], and iPhone [18] dataset. The 3D metrics are obtained by comparing the back-projected depth and our culled mesh of each frame.
Method Metric AMA [67] BANMo [78]
Avg.
Samba
Swing
eagle1
eagle2
NDR Acc. [cm] $\downarrow$
2.32
2.70
6.51
15.42
6.73
Comp. [cm] $\downarrow$
1.94
2.17
2.58
3.70
2.59
Clip sim. $\uparrow$
86.72
88.03
86.23
81.64
85.65
Ours Acc. [cm] $\downarrow$ 1.98 1.71 2.16 4.43 2.57
Comp. [cm] $\downarrow$ 1.88 1.83 2.07 3.67 2.36
Clip sim. $\uparrow$ 92.48 91.86 90.63 90.28 91.31
Table 2: Per-scene quantitative results on synthetic datasets, including AMA [67] and BANMo [78] datasets. The 3D metrics are obtained by comparing the ground-truth mesh of each frame.
Datasets. We evaluate our method on 9 real-world scenes that are from KillingFusion [59], DeepDeform [8], and iPhone dataset [18]. All videos in those datasets are captured with a consumer-grade RGB-D camera. Object masks are obtained with off-the-shelf object segmentation methods [13, 33] following NDR [9]. Additionally, we also evaluate our methods on 3 synthetic scenes from AMA [67] and BANMo [78] dataset which ground-truth meshes at all time-stamps are available. For AMA dataset [67], we render depth images from the calibrated 8 views aligned with the provided GT RGB images and masks, and select the front-view observations for optimisation. For BANMo dataset [78], we follow a similar protocol except that all the RGB, depth and mask are rendered. Following NDR [9], we also compensate for the object’s rigid motion and relative transformation in the data pre-processing stage. Please refer to our supplementary for more details.
Metrics. We adopt accuracy (acc. [cm]) and completion (comp. [cm]) to evaluate the surface reconstruction quality and CLIP similarity to evaluate the realism of completion. For 9 real-world scenes, we perform mesh culling to evaluate the observed region due to the lack of complete 360° GT meshes. The CLIP similarity is computed between RGB frames and the novel view renderings generated by our optimized model. These renderings follow a 360° trajectory around the object, which is in motion over time. In the case of the 3 synthetic scenes, we evaluate the complete mesh and render RGB images from the other 7 views to obtain CLIP similarity.
Implementation Details. We run MorpheuS on a desktop PC with an Intel Core i7-13700K CPU and NVIDIA RTX 4090 GPU with 24GB memory. The model takes around 2-3 hours to train and requires 10-22GB of memory depending on the the size of the input frames. We employ a coarse-to-fine training strategy that progressively increases the bandwidth of the positional encoding and hash encoding by a modulation term [45, 9, 73]. Additionally, we discover that when the learning rate is small, the learning of the canonical field (with hash encoding) is significantly faster than the deformation field (with positional encoding). This can let the model learn a good initialization of the canonical shape. Thus, we apply a learning rate warm-up strategy. In the warm-up phase, we freeze the deformation field when training the novel view with ${\mathcal{L}}_{\mathrm{S}}$ and set the timestep range of the diffusion prior to be $[0.02,0.5]$ to ensure the model learns a good canonical shape representation. In the second phase, we reduce the timestep range to $[0.02,0.2]$ to allow the model to learn accurate motion. The rendered resolution is doubled to improve the quality of the geometry and appearance completion. Please refer to supplementary material for more details.
Acc. [cm]$\downarrow$ Comp. [cm]$\downarrow$ Clip sim. $\uparrow$
w/o diffusion prior 0.99 0.73 82.68
w/o temporal condition 0.92 0.95 86.01
w/o angle weight 0.91 0.86 86.65
w/o depth 3.24 4.90 85.35
w/o canonical space 0.90 1.05 86.14
Full model 0.88 0.78 86.77
Table 3: Ablation studies on real-world datasets: a) w/o diffusion prior: optimize on real-view only; b) w/o depth: optimize with rgb only; c) w/o temporal condition: condition on first frame d) w/o angle weight: using original $w(i)$. e) w/o canonical space: modeling the scene in observation space via a quad-cube, which decomposes the 4D scene into 4 Hash feature grids.
4.2 Evaluation
Tab. 1 shows the per-scene quantitative results across three real-world datasets. For a more comprehensive evaluation, we extend our evaluations to include 4 synthetic scenes in Tab. 2. Both results show that our model demonstrates superior accuracy and completion in surface reconstruction, outperforming NDR which exhibits spurious surface completion due to the continuity MLPs (See Teddy in Fig. 3). Additionally, our method achieves a significantly better CLIP similarity in comparison to NDR, thanks to the effective knowledge distillation from the diffusion prior.
Fig. 3 shows a qualitative comparison of dynamic 360° reconstruction, highlighting the ability of diffusion priors to eliminate spurious surface completion and achieve photo-realistic reconstruction in unobserved regions. Notably, our model exhibits robustness in completing unobserved regions even in the presence of significant motion changes, as illustrated in the last row of the Teddy sequence.
4.3 Analysis
We further analyze different components of our proposed method by conducting a set of ablation experiments on the 9 real-world scenes with quantitative results shown in Tab. 3 and qualitative demonstrations in Fig. 4-7.
Temporal View-dependent SDS. We start by analyzing the diffusion prior. As in Tab. 3, our full model (last row) exhibits a significant improvement in CLIP similarity in 360° compared to the variant without diffusion priors (first row). Notably, the results also highlight the diffusion priors as a regularization for surface reconstruction in the observed region, leading to the improvement in reconstruction accuracy (0.88 v.s. 0.99) by effectively eliminating spurious surfaces.
We further analyze a variant without the temporal condition (second row), where the first frame is used as the condition for the diffusion priors all the time. Results show a degradation in both 3D metrics and CLIP similarity, demonstrating the importance of accurate temporal conditions for estimating the scene geometry when the target object is in motion over time. However, as opposed to using a given frame as a direct data-term supervision, the use of SDS with slightly inconsistent view conditions could still achieve realistic scene completion without drastically affecting real-view reconstruction results.
Additionally, we explore a variant w/o angle weight, results in Tab. 3 show angle weight can effectively improve the robustness of motion capture by weighting the ${\mathcal{L}}_{\mathrm{S}}$ based on the relative angle between real view and sampled view.
Color view 1 Color view 2 Mesh view 1 Mesh view 2
w/o geo. init. Refer to caption Refer to caption Refer to caption Refer to caption
w/ geo. init. Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Ablation study on geometric initialization. Geometric initialization with fixed blob function [47] fails here.
w/o ${\mathcal{L}}_{\mathrm{smooth}}$ w/ ${\mathcal{L}}_{\mathrm{smooth}}$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Ablation study on surface normal smoothness ${\mathcal{L}}_{\mathrm{smooth}}$.
Reference Ours Ours (w/o depth) Magic123 (Coarse) Magic123 (Fine)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Ablation study: RGB+D vs. RGB-only. We train Magic123 [49] on a single frame for reference. Magic123 (Coarse) is the first stage of training using NeRF. Magic123 (fine) is the second stage of training using DMTet [55] that directly optimizes surface mesh. Due to the scale ambiguity, RGB-based methods fail to capture the accurate pose of the human arm.
Geometric Initialization. As in Fig. 4, without the geometric initialization of the canonical SDF field, using SDS loss may not let the model converge to a coherent mesh. Note that here using fixed blob function [47] fails to converge.
Normal Smoothness. Fig. 5 shows that applying normal smoothness loss ${\mathcal{L}}_{\mathrm{smooth}}$ to near-surface points in observation space leads to smoother reconstruction and better motion capture (e.g. hand).
Effect of Depth Measurement. As in Tab. 3, despite we could achieve competitive CLIP similarity without depth measurement, it fails to capture accurate motion and geometry from the monocular RGB video. We further illustrate the pitfalls of the RGB-based methods in Fig. 6. Magic123 [49] is trained on a single frame with a monocular depth prior as a reference. Due to the scale ambiguity, it is very challenging to recover the accurate geometry (e.g., the human arm) for RGB-based methods.
Canonical Space Regularization. We also evaluate the effect of different strategies when applying our regularization losses in Eq. 13. Apart from our proposed canonical space regularization (Eq. 12), we also include two variants obs. perturb. and cano. perturb. both of which involve the deformation network. The perturbation vector $\Delta{\mathbf{x}}$ is applied in the observation space and the canonical space respectively. Results in Fig. 7 reveal that both obs. perturb. and cano. perturb. fails to complete the tail of the cat, whereas our proposed canonical space regularization could enforce the 3D shape in hyper-canonical space to be consistent, leading to better completion of unobserved regions.
Reference obs. perturb. cano. perturb. w/o reg. w/ cano reg.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Ablation study on canonical regularization: a) obs. perturb: w/ deformation field; apply perturbation in observation space. b) cano. perturb: w/ deformation field; apply perturbation in canonical space c) w/o reg: w/o per-point normal regularization d) w/o deformation field; apply perturbation in canonical space.
Effect of Representations with a Canonical Shape. To evaluate the impact of the canonical space, we compare our full model to a canonical space-free representation (w/o canonical space). In the latter, the world space is directly modeled as a 4D space-time grid, which is decomposed into a quad-cube representation with 4 3D hash grids. Without the implicit regularization provided by a canonical space, the quad-cube demonstrates worse temporal consistency in both geometry and appearance reconstruction, as reflected in the quantitative results in Tab. 3.
Input frames Reconstruction
Refer to caption Refer to caption Refer to caption
Figure 8: Limitations: Our method can fail in challenging situations such as incomplete views, motion blur, and challenging articulated poses of the target object.
Limitations. While we demonstrate photo-realistic 360° reconstruction, certain limitations are inherent in our current formulation. Specifically, real-world captures often present challenges such as incomplete views, motion blur, and intricate articulated poses of the target object (see Fig. 8). Our model, which relies on the diffusion prior conditioned on pure RGB frames, may fail in these scenarios and result in incoherent reconstruction, missing geometry, etc. In addition, target objects might also have complex motion in the self-occluded region, which is difficult to model due to the lack of motion priors. Introducing better diffusion priors (e.g., diffusion models that are conditioned on RGB-D images, time, etc) and motion priors can be a promising avenue for future research.
5 Conclusion
We presented MorpheuS, a novel framework designed for dynamic 360° surface reconstruction from a casual monocular RGB-D video. Through the effective integration of the diffusion prior with the dynamic scene reconstruction, MorpheuS goes beyond the conventional dynamic reconstruction methods by achieving both photo-realistic completion in unobserved regions and accurate motion and geometry reconstruction in the observed regions.
Acknowledgements. The research presented here has been supported by a sponsored research award from Cisco Research and the UCL Centre for Doctoral Training in Foundational AI under UKRI grant number EP/S021566/1. This project made use of time on Tier 2 HPC facility JADE2, funded by EPSRC (EP/T022205/1).